OpenStudy (anonymous):

Jake is comparing the prices of two mattress cleaning companies. Company A charges $30 per mattress and an additional $13 as service charges. Company B charges $28 per mattress and an additional $15 as service charges.

OpenStudy (anonymous):

Part A: Write equations to represent Company A's and Company B's total mattress cleaning charges for a certain number of mattresses. Define the variable used in the equations. Part B: Which company would charge less for cleaning 4 mattresses? Justify your answer. Part C: How much money is saved by using the services of Company B instead of Company A to clean 7 mattresses?

OpenStudy (phi):

Company A charges $30 per mattress and an additional $13 as service charges. if you have "x" mattresses, how much do they charge ?

OpenStudy (anonymous):

30x

OpenStudy (phi):

and an additional $13 as service charges.

OpenStudy (anonymous):

so 30x + 13

OpenStudy (phi):

yes

OpenStudy (phi):

if we call the total amount charged f(x), you write f(x) = 30x + 13 or, if you call the total amount charged "y", you write y= 30x +13 now for B: Company B charges $28 per mattress and an additional $15 as service charges.

OpenStudy (anonymous):

28x + 15

OpenStudy (phi):

so write it y= 28x+15 where x is the number of mattresses, and y is the total cost

OpenStudy (phi):

That is part A. Part B: Which company would charge less for cleaning 4 mattresses? Justify your answer. any idea how to do this ?

OpenStudy (anonymous):

wait one minute though, this is how i wrote them: Company A charges 30x + 13 for service charges lets call the total amount charged f(x) f(x) = 30x + 13 Company B charges 28x + 15 for service charges lets call the total amount charged y y = 28x + 15

OpenStudy (phi):

ok, but Company A charges 30x + 13 for service charges does not read well. It sounds like the service charge is 30x+13 I would write Company A charges 30x + 13 to clean x mattresses

OpenStudy (anonymous):

ok i fixed it, ok now for part b?

OpenStudy (phi):

yes. replace "x" with 4 in both equations and see which gives the smaller number (cheaper price)

OpenStudy (anonymous):

f(4) = 30 * 4 + 13 or 30 * 4 + 13

OpenStudy (phi):

you can simplify that to a single number

OpenStudy (anonymous):

f(4) = 30 * 4 + 13 = 133

OpenStudy (phi):

and how much does the other company charge to clean 4 mattresses?

OpenStudy (anonymous):

y = 28 * 4 + 15 = 127

OpenStudy (phi):

now you know the answer to part B, and why

OpenStudy (anonymous):

ok and for part c I do the same thing but with seven instead of 4